### v2

@@ 4 Dilute emulsions: small deformation theory and constitutive models
In this section, we summarize key features of theoretical and numerical investigations of single-drop dynamics and rheology of dilute emulsions by including three cases: clean drops, surfactant-covered drops, and drops with slip at interfaces. We revisit significant theoretical advances made analytically in the two asymptotic limit of small or large droplet deformations in viscometric flows [30, 105]. We mention numerical studies used for bridging the gap between the two asymptotic limits for clean drops [106, 107, 68], surfactant-covered droplets [59, 108, 89, 109, 110, 111, 112, 113, 114], and drops with viscous interfaces [94, 95, 115, 116, 96, 117]. The approaches discussed here form the starting point for investigations on emulsions containing interfaces with non-Newtonian interfacial rheology. For example, proteins or particles as emulsifiers lead to interfacial viscoelasticity or interfacial yield stress and presence of lipid membranes and protein gel networks at interface creates bending modulus, manifested in suspensions of vesicles and cells including blood. We recommend recent reviews and papers for discussions of emulsions containing complex interfaces that exhibit non-Newtonian interfacial rheology.
4.1 Small deformation theories
Taylor’s deformation parameter In 1932, Taylor generalized Einstein’s formula for viscosity a dilute suspension of hard spheres to derive an expression for the viscosity of dilute emulsions in the limit of low $Ca$, clean interface, and for cases with Newtonian dispersed and suspending fluids. Taylor [6, 118] was the first to theoretically and experimentally study the deformation of a neutrally buoyant viscous drop in response to imposed shear or extensional flows, and describe how bulk rheology is informed by drop deformation and orientation at the microscopic scale. For a weakly perturbed spherical drop, the shape change can be measured using a scalar quantity called Taylor’s deformation parameter defined as
@@ 4.1 Small deformation theories
where $L$ and $B$ are the major and minor axes of the ellipsoid projected onto the velocity-shear rate plane, as shown in Figure 2. In flows with a rotational component of velocity including viscometric shear flows, the ellipsoid (projection of the deformed drop) orients forming inclination angle, $\theta$ measured between the major axis of deformation and the flow direction. For large deformations, especially those encountered in response to extensional flows, $L/B$ is usually used instead of $D_{T}$ [105].
@@
In dilute emulsions, the flow-induced droplet dynamics depend on the physicochemical properties of the two liquids (density and viscosity), composition-dependent properties of the interface (interfacial tension, interfacial rheology, and surface forces), and the strength and type of imposed flow fields (shear and extensional). Qualitatively, the extend of drop deformation and orientation for clean droplets is influenced by an interplay of viscous and capillary stresses dependent on $Ca$, defined appropriately by accounting for interfacial tension, deformation rate, and viscosity ratio $\lambda$ ranging 0 to $\infty$. Droplets may attain steady shapes or undergo transient flow-induced deformation, possibly leading to interfacial instabilities and breakup (e.g., tip-streaming, burst, and thread breakup by Rayleigh instabilities) [27, 119, 30, 105]. We recommend the classical papers by Leal’s group for a comprehensive survey of clean droplet dynamics in unbounded shear and extensional flows [29, 28], and direct interested readers to Guido’s review on droplet deformation in confined flows and viscoelastic fluids [77].
@@
Clean droplet dynamics in unbounded shear flows In weak flows, $Ca\ll 1$, steady shapes are nearly spherical and the inclination angle $\theta\sim 45^{\circ}$, to leading order in $Ca$, as sketched in Figure 2. At higher flow strengths, for a given $\lambda$, droplet shapes become more elongated as $Ca$ increases and the major axis of deformation aligns with the flow direction as the droplet rotates in response to the local vorticity of the flow. In this limit, drops with viscosities below a critical value $\lambda_{c}\sim 4$ may undergo breakup at a critical flow strength $Ca_{c}$, whereas high-viscosity drops remain stable for $\lambda>\lambda_{c}$, for arbitrary $Ca$ [30, 105]. For example, clean droplets with the same viscosity as the suspending medium undergo breakup at a critical value $Ca_{c}\approx 0.43$ [120].
@@
Experiments by Mason and coworkers [121, 27] characterized the drop deformation and breakup modes of clean droplets in shear flow for $\lambda<\lambda_{c}$ and some cases reproduced are in Fig.3(a). The breakup modes depend on a balance between the rate of increase of capillary number up to and across $Ca_{c}$ and droplet relaxation time. For $\lambda<0.2$ and high $dCa/dt$ rates, the droplets experience tip-streaming breakup mode; whereas for low enough $dCa/dt$, tip-streaming breakup may be suppressed and the droplet deforms into a thin-liquid thread and breakup into smaller droplets by Rayleigh instability. However, numerical and experimental results in extensional flows support the assumption that tip-streaming instabilities occur only in the presence of surfactants [29, 122, 109, 96]. Theoretical and numerical analysis on tip-streaming breakup instability remains an active area of research.
@@
In weak extensional flows, clean droplets attain a stable, stationary shape for all $\lambda$, where the droplet principal axis of deformation is aligned with the flow direction of maximum extension, as illustrated in Fig. 3(b). Here, the transient approach to steady shapes is monotonic, since the flow is free of vorticity. For $Ca=O(1)$, two main regimes of droplet steady deformation are of interest: (i) nearly ellipsoidal shapes are observed for moderate and large $\lambda$, (ii) for $\lambda\lesssim 0.1$, droplets deform into shapes with nearly-pointed ends. For larger values of $Ca$, high-viscosity drops deform into slender threads that eventually breakup into smaller droplets. Low-viscosity drops are able to sustain highly elongated shapes for even larger flow strengths, but will breakup into small droplets via Rayleigh-Plateau instability if $Ca\gg Ca_{c}$. Drop relaxation after the flow field is switched off may also lead to drop breakup into a chain of droplets of uniform size if the droplet initial shape is sufficiently elongated by the flow.
@@
Figure 3: Schematic diagram of droplet deformation in shear and extensional flows. Image adapted from Ref. [121] for shear flow experiments (a) and from Fig. 9 in Ref. [108] for numerical results in extensional flows (b). The two sets in extensional flows depict snapshots of drop relaxation of clean and surfactant-covered droplets at different dimensionless times as indicated. Details on the experimental data sets in part (a) are listed in Appendix B.
@@
The presence of surface inclusions (e.g., surfactant molecules, proteins, lipids) alters the classical dynamics of transient and steady shapes of clean droplets [18]. For surfactant-covered drops, deviations from the clean droplet deformation are governed by a balance among (i) interface convection of surfactants towards regions of high curvature and stagnation points lowering surface tension locally, (ii) local surfactant dilution due to drop deformation and creation of surface area, and (iii) diffusion of surfactant which tends to homogenize the surfactant distribution along the interface. Gradients in surface tension induce Marangoni stresses which act against surface deformation [19, 59, 55]. The critical $Ca_{c}$ for the onset of unsteady deformation and breakup is usually larger compared to clean droplet results, but it can be smaller depending on flow strength and on the local vorticity of the flow [69]. Figure 3(b) shows the relaxation of clean and surfactant-covered droplets at different times after being initially deformed by an extensional flow. Surfactant redistribution along the droplet surface stabilizes the shape against transient configurations that may lead to droplet breakup.
@@
The qualitative behavior of droplets with viscous interfaces in linear flows introduces an additional surface viscous stress to the force balance Eq. (6), where droplet shape and rheology depend on flow type and emulsion’s composition, for example, the relative contribution of shear and dilatational surface viscosities and their relation to surface pressure and surface tension [115, 95, 55, 117].
@@ 4.2 Constitutive models for dilute emulsions
In the limit when a suspended neutrally bouyant, clean droplet deviates from sphericity only slightly, the droplet surface is given by [9, 8, 78]
where $\epsilon\ll 1$ is a perturbation parameter, $\mathbf{A}$ is the shape distortion tensor, $a$ is the radius of the undeformed, spherical droplet, and $r=(\mathbf{x}\cdot\mathbf{x})^{1/2}$. The rate of change of the droplet shape depends on the kinematics of the imposed flow $\mathbf{u}^{\infty}=(\mathbf{E}+\mathbf{W})\cdot\mathbf{x}$, where $\mathbf{E}$ and $\mathbf{W}$ are the imposed-flow rate-of-strain and vorticity tensors, respectively. The distortion tensor $\mathbf{A}$ can be used to calculate Taylor deformation parameter, inclination angle in shear flows, and define rheological material properties of the fluid. The evolution of the distortion tensor in a reference frame that translates and rotates with the droplet is [30, 74]
@@
$-c_{1}(\lambda)\epsilon\mathbf{A}+O(\epsilon Ca,\epsilon^{2})\,,$ (27)
@@
where $c_{0}(\lambda)=5/(2\lambda+3)$, $c_{1}(\lambda)=40(\lambda+1)/[(19\lambda+16)(2\lambda+3)]$. Dimensionless quantities are defined as $\bar{t}=t\,/(\mu a/\sigma)$, $\bar{\mathbf{E}}=\mathbf{E}/\dot{\gamma}$,$\bar{\mathbf{W}}=\mathbf{W}/\dot{\gamma}$, and $|\mathbf{A}|=1$. A derivation of Eq. (4.2) is shown in Appendix A for completeness. The equation captures how the rate of change of $\mathbf{A}$ is contributed by two competing terms. The first term distorts away from spherical shape and is linearly dependent on the rate of strain, whereas the second term restores unperturbed shape and depends on $\mathbf{A}$. The neglected terms of $O(\epsilon^{2})$ correspond to harmonics higher than second, wheras of $O(\epsilon Ca)$ arise from the straining flow acting on the distorted shape [30].
@@
The form of Eq. (4.2) reveals two small deformation regimes: (i) for weak flows (i.e., $\epsilon\sim Ca\ll 1$ and $\lambda=O(1)$, the distortion is limited by strong interfacial tension effect, and (ii) large-$\lambda$ and arbitrary $Ca$ but not too large for flows with sufficient vorticity where $\epsilon\sim\lambda^{-1}\ll 1$. For a given flow type and small parameter $\epsilon$, Eq. (4.2) is solved for the distortion tensor $\mathbf{A}$. Here, we summarize up to second-order deformation theories for clean droplet deformation and rheology in viscometric flows and include results for surfactant-covered drops, interfacially viscous drops, and drops with interfacial slip conditions.
@@
Clean droplets in shear flows For a clean droplet in weak shear flows where $\epsilon=Ca\ll 1$ and $\lambda=O(1)$ [118], the deformation parameter shows a linear dependence on $Ca$ to leading order
$D_{T}=\frac{19\lambda+16}{16\lambda+16}Ca+O(Ca^{2})$ (28)
and the inclination angle is [53]
$\theta=\frac{\pi}{4}-\frac{(19\lambda+16)(2\lambda+3)}{80(1+\lambda)}Ca+O(Ca^{2})\,.$ (29)
@@
The other limit when $Ca=O(1)$ and $\epsilon=\lambda^{-1}\ll 1$, the leading order solutions for the Taylor deformation parameter and inclination angle are
$D_{T}=\frac{5}{4}\lambda^{-1}+O(\lambda^{-2})\quad\theta=\frac{10}{19}\frac{\lambda^{-1}}{Ca}+O(\lambda^{-2})\,.$ (30)
@@
Higher-order theories have been developed; for detailed derivation and formulas see Refs. [123, 78, 72, 74, 55].
@@
Small deformation theory reveals the characteristic rheological behavior of dilute emulsions. For clean drops in shear flows in the weak flow limit when $\epsilon=Ca\ll 1$ and arbitrary $\lambda$, a second-order deformation analysis yields [26, 58, 72]
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5\lambda+2}{2(\lambda+1)}-D_{0}(\lambda)D_{1}(\lambda)Ca^{2}+O(Ca^{3})\,,$ (31)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=10D_{0}(\lambda)^{2}Ca\,,$ (32)
$\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}-Ca\,D_{0}(\lambda)\frac{3(12+9(1+\lambda)-25(1+\lambda)^{2})}{28(1+\lambda)^{2}}$ (33)
@@
where the coefficients $D_{0}$ and $D_{1}$ are listed in Appendix A.1. Equations (31)-(33) reveal the characteristic shear-thinning behavior of emulsion flows with finite positive and negative first and second normal stress differences.

### v3

@@ 4 Dilute emulsions: small deformation theory and constitutive models
In this section, we summarize key features of theoretical and numerical investigations of single-drop dynamics and rheology of dilute emulsions by including three cases: clean drops, surfactant-covered drops, and drops with slip at interfaces. We revisit significant theoretical advances made analytically in the two asymptotic limits of small or large droplet deformations in viscometric flows [32, 120]. We mention numerical studies used for bridging the gap between the two asymptotic limits for clean drops [121, 122, 73], surfactant-covered droplets [64, 123, 99, 124, 125, 126, 127, 128, 129], and drops with viscous interfaces [107, 108, 130, 131, 109, 132]. The approaches discussed here form the starting point for investigations on emulsions containing interfaces with non-Newtonian interfacial rheology, or composed of dispersed or suspending fluid with non-Newtonian rheology response. For example, proteins or particles as emulsifiers lead to interfacial viscoelasticity [133] or interfacial yield stress, and the presence of lipid membranes and protein gel networks at interface create bending and elastic moduli manifested in suspensions of vesicles and cells including blood. We recommend recent reviews and papers for discussions of emulsions containing complex interfaces that exhibit non-Newtonian interfacial rheology or emulsions formed by using a non-Newtonian dispersed or suspending fluid [45, 1, 9].
4.1 Small deformation of drops in shear and extensional flows
Taylor’s deformation parameter In 1932, Taylor generalized Einstein’s formula for viscosity of a dilute suspension of hard spheres to derive an expression for the viscosity of dilute emulsions in the limit of low $Ca$, clean interface droplets, and for cases with Newtonian dispersed and suspending fluids. The expression for relative viscosity, $\eta_{r}$=($\eta$/$\mu$) in the limit of low shear rate (or low Ca) is given by
$\eta_{r}=1+\frac{5}{2}\phi\frac{(\lambda+2/5)}{(\lambda+1)}\ =1+\frac{5}{2}\phi g_{T}(\lambda)\,,$ (27)
where $g_{T}(\lambda)={(\lambda+2/5)}/{(\lambda+1)}$ is Taylor’s viscosity factor. Recalling that the specific viscosity $\eta_{sp}$=($\eta_{r}$-1) equals the ratio of the contribution of dispersed and suspending phases to viscosity we deduce, from Taylor’s expression, an alternative form for specific viscosity of emulsions,
$\eta_{sp}/{\phi}={\mathbf{\Sigma}^{p}_{12}}/{\mu{\dot{\gamma}}}=\frac{5}{2}{g_{T}(\lambda)}\,.$ (28)
In the limit of large $\lambda$, Taylor’s viscosity factor goes to unity or $g_{T}(\lambda)=1$ recovering Einstein’s formula for suspensions. Upon defining the specific viscosity as $\eta_{sp}$=($\eta_{r}$-1), Eq. (27) yields $\eta_{sp}=(5/2)\,\phi{g_{T}(\lambda})$. In the limit of vanishingly small $\lambda$, the parameter $g_{T}(\lambda)\to 2/5$, implying the relative viscosity of bubbly fluid is just $\eta_{r}$ = 1+$\phi$ and the specific viscosity of bubbly fluid is $\eta_{sp}$ = $\phi$.
Taylor [11, 134] was the first to theoretically and experimentally study the deformation of a neutrally buoyant viscous drop in response to imposed shear or extensional flows, and describe how bulk rheology is informed by drop deformation and orientation at the microscopic scale. For a weakly perturbed spherical drop, the shape change can be measured using a scalar quantity called Taylor’s deformation parameter defined as
@@ 4.1 Small deformation of drops in shear and extensional flows
where $L$ and $B$ are the major and minor axes of the ellipsoid projected onto the velocity-shear rate plane, as shown in Figure 2. For large deformations, especially those encountered in response to extensional flows, $L/B$ is usually used instead of $D_{T}$ [120]. According to Taylor, at steady state, the small deformation of a clean droplet in response to weak shear flows [134], exhibits
$D_{T}=\frac{(19\lambda+16)}{(16\lambda+16)}Ca+O(Ca^{2})=d_{T}(\lambda)Ca+O(Ca^{2})\,.$ (30)
Here, for brevity’s sake, we define the viscosity ratio dependent prefactor as the Taylor’s deformation prefactor ${d_{T}}={(19\lambda+16)}/{(16\lambda+16)}$.
Inclination or orientation of deformed drops In flows with a rotational component of velocity including viscometric shear flows, the ellipsoidal drop (or ellipsoid projection of the deformed drop) orients. An inclination angle, $\theta$ can be measured between the major axis of deformation and the flow direction, as shown in Figure 2. Chaffey and Brenner [58] computed the inclination angle exhibited by perturbed drops in weak flows by carrying out perturbation analysis up to second order in $Ca$, leading to the following expression
$\theta=\frac{\pi}{4}-\frac{(2\lambda+3)}{5}\frac{(19\lambda+16)}{(16\lambda+16)}Ca+O(Ca^{2})\,,$ (31)
or
$\theta=\frac{\pi}{4}-\frac{d_{T}}{c_{0}}Ca+O(Ca^{2})\,,$ (32)
@@
where $c_{0}(\lambda)=5/(2\lambda+3)$. In dilute emulsions, the flow-induced droplet dynamics depend on the physicochemical properties of the two liquids (density and viscosity), composition-dependent properties of the interface (interfacial tension, interfacial rheology, and surface forces), and the strength and type of imposed flow fields (shear and extensional). Qualitatively, the extend of drop deformation and orientation for clean droplets is influenced by an interplay of viscous and capillary stresses dependent on $Ca$, appropriately defined by accounting for interfacial tension, deformation rate, and viscosity ratio $\lambda$ ranging 0 to $\infty$.
@@
Clean droplet dynamics In weak flows, $Ca\ll 1$, steady shapes are nearly spherical and the inclination angle $\theta\sim 45^{\circ}$, to leading order in $Ca$, as sketched in Figure 2 and first analyzed and visualized by Taylor [134]. For the two extreme values of $\lambda$ i.e., 0 and $\infty$, the deformation prefactor, ${d_{T}}$ ranges between 1 and 1.187, implying that in weak flows, drop deformation parameter ${D_{T}}$ is linearly dependent on $Ca$. The inclination, $\theta$ according to Chaffey and Brenner equation (32) shows a linear dependence on $Ca$. Visualizing drop deformation under mild flow provides a means of measuring interfacial tension, even when the interfacial tension is extremely small, for example, in water-in-water emulsions or in coacervates. Alternatively, the relaxation of a perturbed drop to its unperturbed state after cessation of flow can be used to measure shape relaxation time and interfacial tension. At higher flow strengths, for a given $\lambda$, droplet shapes become more elongated as $Ca$ increases and the major axis of deformation aligns with the flow direction as the droplet rotates in response to the local vorticity of the flow. In this limit, drops with viscosities below a critical value $\lambda_{c}\sim 4$ may undergo breakup at a critical flow strength $Ca_{c}$, whereas high-viscosity drops remain stable for $\lambda>\lambda_{c}$, for arbitrary $Ca$ [32, 120]. For example, clean droplets with the same viscosity as the suspending medium undergo breakup at a critical value $Ca_{c}\approx 0.43$ [135].
@@
Experiments by Mason and coworkers [136, 29] characterized the transient and equilibrium drop shapes for $\lambda<\lambda_{c}$ and observed breakup modes for clean droplets as illustrated by cases reproduced in Fig.3(a). Breakup modes were observed to depend on a balance between the rate of increase of capillary number up to and across $Ca_{c}$ and the shape relaxation time. For $\lambda<0.2$ and high $Ca$ rates, the droplets experience tip-streaming breakup mode; whereas for low enough $Ca$ rates, tip-streaming breakup may be suppressed and the droplet deforms into a thin-liquid thread and breakup into smaller droplets by Rayleigh instability. However, numerical and experimental results in extensional flows support the assumption that tip-streaming instabilities occur only in the presence of surfactants [31, 137, 124, 109]. Theoretical and numerical analysis on tip-streaming breakup instability remains an active area of research.
@@
In weak extensional flows, clean droplets attain a stable, stationary shape for all $\lambda$, where the droplet principal axis of deformation is aligned with the flow direction of maximum extension, as illustrated in Fig. 3(b) adapted from Milliken et al. [123]. Here, the transient approach to steady shapes is monotonic, since the flow is vorticity-free. For $Ca=O(1)$, two main regimes of droplet steady deformation are of interest: (i) nearly ellipsoidal shapes are observed for moderate and large $\lambda$, (ii) for $\lambda\lesssim 0.1$, droplets deform into shapes with nearly-pointed ends. For larger values of $Ca$, high-viscosity drops deform into slender threads that eventually breakup into smaller droplets. Low-viscosity drops are able to sustain highly elongated shapes for even larger flow strengths, but will breakup into small droplets via Rayleigh-Plateau instability if $Ca\gg Ca_{c}$. Drop relaxation after the flow field is switched off may also lead to drop breakup into a chain of droplets of uniform size if the droplet initial shape is sufficiently elongated by the flow.
@@
Figure 3: Schematic diagram of droplet deformation in shear and extensional flows. Image adapted from Ref. [136] for shear flow experiments (a) and from Fig. 9 in Ref. [123] for numerical results in extensional flows (b). The two sets in extensional flows depict snapshots of drop relaxation of clean and surfactant-covered droplets at different dimensionless times as indicated. Details on the experimental data sets in part (a) are listed in Appendix C.
@@
Deformation of coated droplets The presence of surface inclusions (e.g., surfactant molecules, proteins, lipids) alters the classical dynamics of transient and steady shapes of clean droplets [20]. For surfactant-covered drops, deviations from the clean droplet deformation are governed by a balance among (i) interface convection of surfactants towards regions of high curvature and stagnation points lowering surface tension locally, (ii) local surfactant dilution due to drop deformation and creation of surface area, and (iii) diffusion of surfactant which tends to homogenize the surfactant distribution along the interface. Gradients in surface tension induce Marangoni stresses which act against surface deformation [21, 64, 60]. The critical $Ca_{c}$ for the onset of unsteady deformation and breakup is usually larger compared to clean droplet results, but it can be smaller depending on flow strength and on the local vorticity of the flow [74].
@@
Figure 3(b) shows the relaxation of clean and surfactant-covered droplets at different times after being initially deformed by an extensional flow. Surfactant redistribution along the droplet surface stabilizes the shape against transient configurations that may lead to droplet breakup. The qualitative behavior of droplets with viscous interfaces in linear flows introduces an additional surface viscous stress to the force balance Eq. (6), where droplet shape and rheology depend on flow type and emulsion’s composition, for example, the relative contribution of shear and dilatational surface viscosities and their relation to surface pressure and surface tension [130, 108, 60, 132]. Thus, droplets may attain steady shapes or undergo transient flow-induced deformation, possibly leading to interfacial instabilities and breakup (e.g., tip-streaming, burst, and thread breakup by Rayleigh instabilities) [29, 138, 32, 120].
Deformation a drop in an emulsion with non-Newtonian component(s) The deformation of drops in emulsions containing one non-Newtonian phase have received more attention than cases with both non-Newtonian phases. Both experiments, analytical theory and numerical simulations show that viscoelasticity changes both the drop deformation and ${Ca_{c}}$ above which elongated drops breakup. A detailed understanding of interplay of capillarity and viscoelasticity is absent even for drop coalescence and breakup as it requires examination and understanding of the influence of rate-dependent shear viscosity, transient and steady extensional viscosity, normal stress differences, and for polymeric fluids, finite extensibility and non-Hookean elasticity [27, 139, 140, 141, 142, 143, 45, 116]. The experimentalists often utilize constant-viscosity elastic fluids called Boger fluids to isolate the effect of elasticity without interference from rate-dependent viscosity behavior. The constant viscosity elastic fluids can be considered analogs of Oldroyd-B fluids especially while modeling shear rheology response, and for the small deformation case, the simpler second-order fluid model provides a reasonable starting point [116, 9]. However, these models are suitable for comparisons with experiments only for slow flows, but the practical problems of dilute and nondilute emulsions often require an understanding of response to strong flows and coupling between rate-dependent viscous and nonlinear viscoelastic effects [116, 9, 45].
We recommend the classical papers by Leal’s group for a comprehensive survey of clean droplet dynamics in unbounded shear and extensional flows [31, 30], including studies when drop or suspending fluid is non-Newtonian [144, 142, 143], and direct the interested readers to Guido’s review on droplet deformation in confined flows and viscoelastic fluids [45]. Numerical methods used for visualization of drop deformation in nondilute emulsions for different choices of $\lambda$, flow types, $Ca$, interfacial properties, and Newtonian or viscoelastic fluids are summarized in a later section. Next we describe how drop deformation, orientation and shape relaxation provide ingredients for deriving or prescribing constitutive models for emulsions by incorporating the role of interfacial and bulk properties of dispersed and suspending fluids.
@@ 4.2 Constitutive models for dilute emulsions
Evolution of the shape distortion tensor In the limit when a suspended neutrally buoyant, clean droplet of undeformed, spherical radius $a$ deviates from sphericity to slight ellipticity, the perturbed drop shape is described by [15, 14, 88]
$S(t)=r(t)-a\left(1+\epsilon\frac{\mathbf{x}\cdot\mathbf{A}(t)\cdot\mathbf{x}}{r^{2}}\right)+O(\epsilon^{2})=0\,,$ (33)
where $\epsilon\ll 1$ is a perturbation parameter, and $r=(\mathbf{x}\cdot\mathbf{x})^{1/2}$. The shape distortion tensor, $\mathbf{A}$, measures the droplet deformation embedded in the current configuration relative to a reference configuration. The components of $\mathbf{A}$, a second-order tensor, are determined in terms of a second-order deformation gradient tensor, $\mathbf{F}$, that maps the deformation of the material lines from reference to current configuration [145]. Please refer to Appendix A for a more detailed description.
The rate of change in the droplet shape depends on the kinematics of the imposed flow $\mathbf{u}^{\infty}=(\mathbf{E}+\mathbf{W})\cdot\mathbf{x}$, and thus on the rate-of-strain tensor $\mathbf{E}$ and vorticity tensor $\mathbf{W}$. The distortion tensor $\mathbf{A}$ can be used to calculate the Taylor deformation parameter ${D_{T}}$, orientation (inclination angle in shear flows), and define rheological material properties of the emulsion. The evolution of the distortion tensor in a reference frame that translates and rotates with the droplet [32, 78] is captured by the following expression
@@
$-c_{1}(\lambda)\epsilon\mathbf{A}+O(\epsilon Ca,\epsilon^{2})\,.$ (34)
@@
Here the two coefficients $c_{0}(\lambda)=5/(2\lambda+3)$ and $c_{1}(\lambda)=40(\lambda+1)/[(19\lambda+16)(2\lambda+3)]$ depend primarily on viscosity ratio, $\lambda$. The coefficient $c_{0}(\lambda)$ appeared in the definition of the deformation parameter, $D_{T}$ in Eq. (32). The dimensionless quantities are defined for time, strain rate tensor and vorticity tensor, respectively as $\bar{t}=t\,/(\mu a/\sigma)$, $\bar{\mathbf{E}}=\mathbf{E}/\dot{\gamma}$,$\bar{\mathbf{W}}=\mathbf{W}/\dot{\gamma}$, and $|\mathbf{A}|=1$. A derivation of Eq. (4.2) is included in the Appendix A for completeness. The right-hand side of Eq. (4.2) captures how the rate of change of $\mathbf{A}$ is contributed by two competing terms. The first term distorts away from spherical shape and is linearly dependent on the rate of strain, whereas the second term restores unperturbed shape and depends on $\mathbf{A}$. The neglected terms of $O(\epsilon^{2})$ correspond to harmonics higher than second, whereas terms of $O(\epsilon Ca)$ arise from the straining flow acting on the distorted shape [32]. It is possible to rewrite $c_{1}(\lambda)=1/[2{c_{0}}{d_{T}}]$ and reframe the Eq. (4.2) to appear as:
$\epsilon\frac{\partial\mathbf{A}}{\partial\bar{t}}-Ca\bar{\mathbf{W}}\cdot\epsilon\mathbf{A}+\epsilon Ca\mathbf{A}\cdot\bar{\mathbf{W}}=Ca\,c_{0}\bar{\mathbf{E}}$
$-\frac{1}{2c_{0}d_{T}}\epsilon\mathbf{A}+O(\epsilon Ca,\epsilon^{2})\,.$ (35)
@@
The form of Eq. (4.2) or Eq. (4.2) reveals two small deformation regimes: (i) for weak flows (i.e., $\epsilon\sim Ca\ll 1$ and $\lambda=O(1)$), the distortion is limited by strong interfacial tension effect, and (ii) large-$\lambda$ and arbitrary $Ca$ but not too large for flows with sufficient vorticity where $\epsilon\sim\lambda^{-1}\ll 1$. For a given flow type and small parameter $\epsilon$, Eq. (4.2) is solved for the distortion tensor $\mathbf{A}$. Here, we summarize up to second-order deformation theories for clean droplet deformation and rheology in viscometric flows and include the results for surfactant-covered drops, interfacially viscous drops, and drops with interfacial slip conditions. The interfacial slip case is key to understanding the rheology of polymer blends as emulsions, which are formed by phase separation.
@@
Clean droplets in shear flows For a clean droplet in weak shear flows where $\epsilon=Ca\ll 1$ and $\lambda=O(1)$, the deformation parameter ${D_{T}}$ derived by Taylor [134] and the inclination angle $\theta$ as demonstrated by Chaffey and Brenner [58] shows a linear dependence on $Ca$. These expressions discussed in the previous section and given by Eqs. (30) and (32), respectively, are reproduced here for clarity of presentation,
$D_{T}=\frac{19\lambda+16}{16\lambda+16}Ca+O(Ca^{2})={d_{T}}Ca+O(Ca^{2})\,,$ (36)
$\theta=\frac{\pi}{4}-\frac{d_{T}}{c_{0}}Ca+O(Ca^{2})\,.$ (37)
@@
In the other limit when $Ca=O(1)$ and $\epsilon=\lambda^{-1}\ll 1$, the leading order solutions for the Taylor deformation parameter and inclination angle are
$D_{T}=\frac{5}{4}\lambda^{-1}+O(\lambda^{-2})\,,\quad\theta=\frac{10}{19}\frac{\lambda^{-1}}{Ca}+O(\lambda^{-2})\,.$ (38)
@@
Higher-order theories have been developed; for detailed derivation and formulas, see Refs. [146, 88, 77, 78, 60].
@@
For clean drops in shear flows in the weak flow limit when $\epsilon=Ca\ll 1$ and arbitrary $\lambda$, a second-order deformation analysis [13, 63, 77] leads to the following equations that describe the characteristic rheological behavior of dilute emulsions
$\frac{\Sigma^{p}_{12}}{\mu\dot{\gamma}}=\frac{5}{2}{g_{T}}-\frac{4}{5}{d_{T}}D_{1}(\lambda)Ca^{2}+O(Ca^{3})\,,$ (39)
$\frac{N^{p}_{1}}{\mu\dot{\gamma}}=\frac{32}{5}{d_{T}}^{2}Ca\,,$ (40)
$\frac{N^{p}_{2}}{\mu\dot{\gamma}}=-\frac{1}{2}\frac{N^{p}_{1}}{\mu\dot{\gamma}}-\frac{4}{5}Ca\,{d_{T}}D_{2}(\lambda)$ (41)
@@
where the coefficients $D_{0}$ and $D_{1}$ are listed in Appendix A.1. Equations (39)-(41) reveal the characteristic shear-thinning behavior of emulsion flows with finite positive and negative first and second normal stress differences, respectively.
Since ${\mathbf{\Sigma}^{p}_{12}}/{\mu{\dot{\gamma}}}=\eta_{sp}/{\phi}$, the explicit use of ${g_{T}}$, ${d_{T}}$ and ${c_{0}}$ in writing the contributions of dispersed phase (or particulates) to shear stress, specific viscosity, and normal stress differences provides two key benefits. The expressions appear more compact and allow for clearer comparisons. Also, from $\lambda$ ranging from zero (bubble) to infinity (rigid particle), ${d_{T}}$ varies from 1 to 1.187, whereas ${g_{T}}={(\lambda+2/5)}/{(\lambda+1)}$ increases from 2/5 to 1.